determine the heat evolution inside a horizontal tube heated at the bottom at a temperature of 40 degree celsius, diameter of the tube is negligible and tube length is 1m, v =1cm/sec


1
Expert's answer
2021-05-28T02:25:49-0400

Heat evolution is the measure of enthalpy "(\triangle h)"

"(\triangle h) =u+PV" , but u, the internal energy is zero

"(\triangle h) =Pv"

"\frac{P}{T}=k \implies P=kT, if \space k=1, P=T=40+273=313K"

"(\triangle h) =313 \times \frac{1cm}{sec} \times\frac{1m}{100cm} \times 1kg=3.13 kg/ms"

"(\triangle h)=31.3 J"